OpenStudy (anonymous):

What are the solutions of the quadratic equation? x2 + 11x = –24 (Multiple Choice) ◘ -3, -8 ◘ -3, 8 ◘ 3, -8 ◘ 3, 8

OpenStudy (anonymous):

1st step?

OpenStudy (anonymous):

That is actually x^2

OpenStudy (anonymous):

try and i will help

OpenStudy (anonymous):

I have no idea where to even start with this, honestly.

OpenStudy (anonymous):

you need to make it look like this x^2 + bx + c =0

OpenStudy (anonymous):

Umm, how do you do that?

OpenStudy (anonymous):

x^2 + 11x = –24 add 24 to both sides

OpenStudy (anonymous):

so its x^2 + 11x + 24 = 0 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now what are the 2 numbers that what numbers can make 24 a*b

OpenStudy (anonymous):

8 and 3

OpenStudy (anonymous):

so its (x + 3)(x + 8) = 0

OpenStudy (anonymous):

yes but -3 and -8 also make 24, why did you choose 8 and 3

OpenStudy (anonymous):

Because they add together to equal 11.. isnt that what you do when factoring?

OpenStudy (anonymous):

correct (x + 3)(x + 8) = 0 now what will make this function true?

OpenStudy (anonymous):

I don't really know what to do from this point.

OpenStudy (anonymous):

(x + 3)(x + 8) = 0 x = -3 (-3 + 3)(x + 8) = 0 0(x + 8) = 0 right?

OpenStudy (anonymous):

Why does x= -3

OpenStudy (anonymous):

to make the function = 0 you need 0*a=a so -3 + 3 = 0 and -8 +8 =0

OpenStudy (anonymous):

okay. . .

OpenStudy (anonymous):

Okay, I understand now, so the solutions are -3 and 8?

OpenStudy (anonymous):

the question is not asking for factoring it asks for the solution (x + 3)(x + 8) = 0 is factoring but how to make the left hand side equal to zero

OpenStudy (anonymous):

why 8?

OpenStudy (anonymous):

Because thats how I saw it in the equation, is it not 8?

OpenStudy (anonymous):

ok lets plug it x= 8 (x + 3)(x + 8) = 0 (8 + 3)(8 + 8) = 0 (11)(16) = 0 ? is it true?

OpenStudy (anonymous):

oh, no.

OpenStudy (anonymous):

So its -8

OpenStudy (anonymous):

you got it -3,-8

OpenStudy (anonymous):

Okay, thank you so much.

OpenStudy (camzzzie):

1. a 2.c 3.a 4.d 5.d 100% qc

OpenStudy (dayianaforever):

yeah she's right just try it

OpenStudy (anonymous):

She is right i just got it right

OpenStudy (anonymous):

100% CORRECT